Kaye spent 1/5 of her money in one
store. In the next store she spent three
times as much as she spent in the first
store, and 60 pesos left. How much
money did she have at the start?

Found 2 solutions by htmentor, greenestamps:
Answer by htmentor(1343)

Let M = the amount of money she started with
In the first store, she spends M/5.
In the second store, she spends 3*(M/5)
Since she is left with 60 dollars, we can write the following equation:
60 = M - M/5 - 3M/5
M/5 = 60 -> M = 300

Answer by greenestamps(13334)



She spent 1/5 of the money at the first store.

At the second store, she spent 3 times as much -- that is, she spent 3/5 of her original amount.

So in the two stores together she spent 4/5 of her money, meaning she has 1/5 of it left.

If 1/5 of her money is 60 pesos, then her original amount was 5*60 pesos = 300 pesos.